68. Object A with mass 8.0 kg travels to the east at 10.0 m/s and object B with mass 3.0 kg travels south at 20.0 m/s. The two objects collide and stick together as shown below. What is the magnitude of the velocity they have after the collision? What is the direction of the velocity they have after the collision? 1- 10 m/s
Answer #1

68) let the x-component of final speed be Vx

and y-component be Vy

conserving the momentum in x-direction we have

m1u1x + m2u2x = (m1 + m2)*Vx

8*10 + 0 = 8Vx + (8 + 3) *Vx

Vx = 7.27 m/s

conserving the momentum in y-direction

0 + 3*-20 = (8+3)Vy

Vy = -5.45 m/s

magnitude = sqrt(7.27^2 + 5.45^2) = 9.08 m/s

direction = tan^-1 ( 5.45 / 7.27) = 36.85o south of east